Home‎ > ‎Crew 135‎ > ‎

205-CommanderReport

Commander Report
02/05/2014
Ondrej Doule
 
Crew 135 is performing well and has great dynamics. The only complaint is about the very dry air
that dropped to 11-15 percent this morning. The air circulation is probably too fast in-out of
habitat requiring more heating than well sealed habitat would require.
 
Crew worked on theoretical research and greehouse research in the morning and performed EVA and in-
hab research in the afternoon. Much experience was acquired during today EVA. The scratched EVA
helmets makes it very difficult to read data on LCD displays and beam aiming. Laser amplifying
glasses helpd with the aiming issue. The team performed fairly complex activities dealing with
precision measurements of the habitat and surrounding clearence.
We acquired lots of data for post processing.
 
We generated first ideas about the MDRS large scale expansion regarding topography.
 
 
Crew engineer is encountering problem with location of lower deck thermostat. It is possible that the temperature was read incorrectly last time – please provide instructions which unit is called Lower Deck Termostat and where this unit is located.

=====

CommanderCheckIn Report
02/05/2014
Ondrej Doule
Crew Physical Health:
All are well.

Eva Depart Time:
16:00
Eva Return Time:
17:30
Narrative Of Field Mission Results:
Most of the measurements of the external strucutres and clearances were acquired. Ergonomics experiment was successfully performed in parallel.
Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Hab Maint:
- Two chargin cables for EVA PLSS backpacks were fixed. Five backpacks can be charged from now on.
- Toilet clogging problem was encountered today, although we carrefully follow the toilet use instructions. Please advice on procedures in case the toilet drainage pipes are not functioning properly.
Rep Sched:
19:00 - 21:00
Tomorrow Plan:
Research in habitat:
EXP01 - Probabilistic modeling ongoing
EXP02 - Greenhouse research is ongoing
EXP03 - Initiation of the structural experiment loding/optical record
EXP05 - Ergonomics experment during EVA 04
EXP06 - MDRS CAD drawings - more time required due to geometrical uncertainities
EXP07 - Nutrition experiment ongoing
- EVA04 - Measuremenets of surrounding terrain in close vicinity of the base while performing ergonomics experiment and performance of the structural/optical experiment.

Inventory:

Support Req:
  Toilet clogging problem was encountered today, although we carrefully follow the toilet use instructions. Please advice on procedures in case the toilet drainage pipes are not functioning properly.
Miscelleneous: